OpenStudy (anonymous):

find the equation of a line containing the given point and having the given slope (4,4) m= 3/8

OpenStudy (anonymous):

hello, ok, do you have the slope-intercept equation? (x,y) = (4,4) is the point and m=3/8

OpenStudy (anonymous):

4 = (3/8) * 4 + b, solve for b. Therefore, b = 4 - (3/8)*4

OpenStudy (anonymous):

b = (8/8)*4 - (3/8)*4 = (5/8)*4 = 5/2 = 2.5

OpenStudy (anonymous):

did this help?

OpenStudy (anonymous):

yes but i am still confused on the fraction part ?

OpenStudy (anonymous):

oh, the 8/8 part?

OpenStudy (anonymous):

that is just for easier operation. you have (3/8)*4, which you could convert to 12/8, then convert 4 to 32/8, right?

OpenStudy (anonymous):

but to get to 32/8, you multiply 4 by (8/8). because both values are 4 * (some fraction)... you have 4* ( 8/8 - 3/8 ) which is 4 *( 5/8 )

OpenStudy (anonymous):

this reduces the actual work you do, because you didn't have to obtain the product of 4 * 8/8